# The asymptotic Plancherel formula and Lusztig’s asymptotic algebra for $\tilde{\mathsf{A}}_{n}$

### v1

@@ Introduction
For general affine (equal parameter) type, Lusztig gave a conjectural description of the asymptotic algebra in terms of the Langlands dual group [20]. In the case of affine type $\mathsf{A}$, this conjecture is equivalent to show that, for all two-sided cells $\Delta_{\lambda}$, the subalgebra $\mathcal{J}_{\lambda}=\mathcal{J}_{\Delta_{\lambda}}$ is isomorphic to the representation ring of a certain connected reductive group $F_{\lambda}$ with Weyl group $G_{\lambda}$. This was first proved by Xi in a remarkable paper [31] in 2002 where he gave the first explicit construction of the asymptotic algebra in affine type $\mathsf{A}$ using the notion of chains and antichains, following work of Shi [28]. Around the same time Bezrukavnikov and Ostrik [4] verified Lusztig’s conjecture in a modified form for general affine type (see also [3]). A further approach to constructing the asymptotic algebra in affine type $\mathsf{A}$ was given recently by Kim and Pylyavskyy [14] using the affine matrix ball construction.
The aim of this paper is to present a new construction of the asymptotic algebra in affine type $\mathsf{A}$ as a matrix algebra with coefficients in the ring of $G_{\lambda}$-symmetric functions which we believe is interesting for three main reasons: (1) our construction is based on the combinatorial notion of alcove paths and the representation ring of $F_{\lambda}$ appears very naturally via the $G_{\lambda}$-Schur symmetric functions, (2) an asymptotic version of the Plancherel Theorem is proved, which endows the asymptotic algebra with an inner product, and (3) there are indications that our methods are adaptable to other affine types, and moreover to the unequal parameter case. For instance, similar methods have been used by the second and fourth authors to construct the asymptotic algebra associated to the lowest two-sided cell in general affine type (for all parameters; see [11, Section 6]) and to construct the full asymptotic algebra for all affine Hecke algebras associated to rank $2$ root systems (for all parameters; see [11, 12]).
@@ 1. $\bullet$
    We develop an asymptotic version of the Opdam’s Plancherel Theorem [24] in affine type $\mathsf{A}$ by showing that the terms in the Plancherel Theorem are in bijection with the set of two-sided cells, and that the coefficients that appear are linked to Lusztig’s $\mathbf{a}$-function (see Section 5.2).
@@ 3. $\bullet$
We now describe in more details the content of the paper. To understand the general philosophy it is helpful to understand the situation for the lowest two-sided cell $\Delta_{0}$ (the minimal cell with respect to the two-sided order $\leq_{LR}$; in this case our method applies to all affine Weyl groups and all choice of parameters). This cell corresponds to the partition $(1^{n+1})$, and contains the longest element $\mathsf{w}_{0}$ of the finite Weyl group $W$. Let $\Gamma_{0}$ be the right cell that contains $\mathsf{w}_{0}$. Then $\Gamma_{0}\cap\Gamma_{0}^{-1}$ is the set of words $t_{\gamma}\mathsf{w}_{0}$ where $\gamma\in P_{+}$ is a dominant weight and $t_{\gamma}\in\widetilde{W}$ is the translation by $\gamma$. Let $\pi_{0}$ be the principal series representation of $\widetilde{H}$ as defined in [11] for example. Then $\pi_{0}$ has a combinatorial description in terms of positively folded alcoves paths, is bounded by $\ell(\mathsf{w}_{0})$ and recognises $\Delta_{0}$ (see [11, Section 6]). Furthermore, the leading matrix $\mathfrak{c}_{0}(t_{\gamma}\mathsf{w}_{0})$ for $\gamma\in P_{+}$ has a unique non-zero coefficient, equal to $\mathfrak{s}_{\gamma}(\zeta)$ (the Schur function of type $W$). In this case, the Satake Isomorphism provides an isomorphism between ${\bf 1}_{0}\widetilde{H}{\bf 1}_{0}$ and the ring of $W$-symmetric functions (see [15, 23]). In this paper we construct analogues of the above results for all two-sided cells.
@@
Before turning our attention to the asymptotic Plancherel Theorem, let us first recall the situation for a finite dimensional Hecke algebra $H_{0}$ of type $\mathsf{A}_{n}$. In this case the canonical trace decomposes as a linear combination of irreducible characters (indexed by partitions) and the coefficients that appear are normalisations of the generic degrees of $H_{0}$. It turns out that the coefficient associated to $\lambda$ has valuation $2\mathbf{a}(\mathsf{w}_{\lambda^{\prime}})$ (see [8]). There is an analogue of this decomposition for affine Hecke algebras in the form of the Plancherel Theorem [24] which expresses the canonical trace as a sum of integrals over families of representations. In this paper we show that, in affine type $\mathsf{A}$, a similar phenomenon as in the finite case happens: the terms in the Plancherel formula are in bijection with the two-sided cells, and the coefficients have valuation equal to $2\mathbf{a}(\Delta_{\lambda})$ (this behaviour was first observed in affine Hecke algebras of types $\tilde{\mathsf{G}}_{2}$ and $\tilde{\mathsf{C}}_{2}$ by the second and fourth authors in [11, 12]). This leads to an asymptotic version of the Plancherel Theorem, which in turn gives rise to an inner product on the ring $\mathfrak{C}_{\lambda}$ of leading matrices.
@@
In Section 5 we recall the Plancherel Theorem for type $\tilde{\mathsf{A}}_{n}$, following [2]. We then develop our asymptotic Plancherel Theorem in Theorem 5.6. This allows us to prove that $\pi_{\lambda}$ recognises $\Delta_{\lambda}$ (Theorem 5.7) and that $\mathcal{J}_{\lambda}\cong\mathfrak{C}_{\lambda}$ (Theorem 5.8).
@@ 1.1 The symmetric group and type $\mathsf{A}_{n}$ root system
The support of a root $\alpha\in\Phi$ is $\mathrm{supp}(\alpha)=\{i\mid c_{i}\neq 0\}$, where $\alpha=\sum_{i=1}^{n}c_{i}\alpha_{i}$. For $J\subseteq\{1,\ldots,n\}$ let $\Phi_{J}=\{\alpha\in\Phi\mid\mathrm{supp}(\alpha)\subseteq J\},$ and for $w\in W$ write $\Phi_{J}(w)=\Phi(w)\cap\Phi_{J}$. If $w=uv$ with $u\in W_{J}$ and $v\in{{}^{J}}W$ then $\Phi(u)=\Phi(w)\cap\Phi_{J}$. In particular, we have ${{}^{J}}W=\{v\in W\mid\Phi_{J}(v)=\emptyset\}$ (see [10, Lemma 1.2]).
@@ 1.2 Partitions and tableaux
Let $\leq$ be the dominance partial order on $\mathcal{P}(n+1)$ given by $\mu\leq\lambda$ if and only if $\mu(i)\leq\lambda(i)$ for all $i\geq 1$ (where, by convention, we set $\lambda_{i}=0$ if $i>r(\lambda)$). Recall that $\mu\leq\lambda$ if and only if $\lambda^{\prime}\leq\mu^{\prime}$.
@@
Then for $j\in J_{\lambda}$ we have $\langle\alpha_{j},\rho_{\lambda}\rangle=1$ (c.f. [10, Lemma 1.4]).
@@ 1.3 The element $\mathsf{u}_{\lambda}$
Let $\mu=\boldsymbol{\mu}(w,\lambda)$. Let $\gamma_{1},\gamma_{2},\ldots,\gamma_{k}$ be the blocks of the $\lambda$-expression for $w$, arranged so that they are in decreasing length. Thus $\mu=(|\gamma_{1}|,\ldots,|\gamma_{k}|)$. Let $\mathbf{t}$ be a (not necessarily column strict) tableau of shape $\mu$ given by entering the elements of $\gamma_{i}$ into the $i$th row of $\mathbf{t}$ in any order. By construction of the $\lambda$-expression, the elements of row $i$ of $\mathbf{t}$ lie in different columns of $\mathbf{t}_{c}(\lambda^{\prime})$, and so by the Dominance Lemma for partitions (see [27, Lemma 2.2.4]) we have $\mu\leq\lambda^{\prime}$ as required. ∎
@@ 1.4 The extended affine Weyl group
For $\alpha\in\Phi$ and $k\in\mathbb{Z}$ let
$H_{\alpha,k}=\{x\in V\mid\langle x,\alpha\rangle=k\},$
and let $s_{\alpha,k}(v)=v-(\langle v,\alpha\rangle-k)\alpha$ be the orthogonal reflection in $H_{\alpha,k}$. The group $W_{\mathrm{aff}}$ is a Coxeter group with generators $S=\{s_{0}\}\cup S_{0}$, where $s_{0}=s_{\varphi,1}$, with $\varphi$ the highest root of $\Phi$. The group $\widetilde{W}$ is not a Coxeter group, however we have $\widetilde{W}=W_{\mathrm{aff}}\rtimes\Sigma$ where $\Sigma=P/Q\cong\mathbb{Z}/(n+1)\mathbb{Z}$. We extend the length function $\ell:W_{\mathrm{aff}}\to\mathbb{N}$ to $\widetilde{W}$ by setting $\ell(w\pi)=\ell(w)$ for all $w\in W_{\mathrm{aff}}$ and $\pi\in\Sigma$. Thus $\Sigma=\{w\in\widetilde{W}\mid\ell(w)=0\}$. Each $\pi\in\Sigma$ induces a permutation of the nodes of the extended Dynkin diagram by $\pi s_{i}\pi^{-1}=s_{\pi(i)}$ for $0\leq i\leq n$. Let $\sigma\in\Sigma$ be the element with $\sigma(i)=i+1$ (with indices read cyclically). By a reduced expression for $w\in\widetilde{W}$ we shall mean a decomposition $w=s_{i_{1}}\cdots s_{i_{\ell}}\pi$ with $\ell=\ell(w)$ and $\pi\in\Sigma$.
@@ 1.5 The extended affine Hecke algebra of type $\tilde{\mathsf{A}}_{n}$
The Bernstein relation (see [19, Proposition 3.6]) is
@@ 1.6 Kazhdan-Lusztig theory
Let $\deg:\mathsf{R}\to\mathbb{Z}\cup\{-\infty\}$ denote degree in $\mathsf{q}$. For example $\deg(\mathsf{q}^{-3})=-3$ and $\deg(0)=-\infty$. Since $h_{x,y,z}$ is bar-invariant we have $\deg(h_{x,y,z})\geq 0$ (provided $h_{x,y,z}\neq 0$), and by [16, §7.2] that $\deg(h_{x,y,z})\leq\ell(\mathsf{w}_{0})$ for all $x,y,z\in\widetilde{W}$.
@@
In particular, note that for $w\in W_{J}$ we have
$T_{w}C_{\mathsf{w}_{J}}=C_{\mathsf{w}_{J}}T_{w}=\mathsf{q}^{\ell(w)}C_{\mathsf{w}_{J}},$
and hence
$C_{\mathsf{w}_{J}}^{2}=\mathsf{q}^{\ell(w_{J})}W_{J}(\mathsf{q}^{-2})C_{\mathsf{w}_{J}}=\mathsf{q}^{-\ell(\mathsf{w}_{J})}W_{J}(\mathsf{q}^{2})C_{\mathsf{w}_{J}}.$
@@
Since $\ell(sw)<\ell(w)$ for all $s\in J$ we have $C_{s}C_{w}=(\mathsf{q}+\mathsf{q}^{-1})C_{w}$ (see [21, Theorem 6.6]), and hence $T_{s}C_{w}=\mathsf{q}C_{w}$. Thus $T_{v}C_{w}=\mathsf{q}^{\ell(v)}C_{w}$ for all $v\in W_{J}$. The result follows by expanding $C_{\mathsf{w}_{J}}$ in the $T_{w}$ basis using (1.3). ∎
@@
Let $\mathcal{D}=\{w\in W\mid\mathbf{a}(w)=-\deg(P_{e,w})\}$. The elements of $\mathcal{D}$ are called distinguished involutions. In [21], Lusztig gave a series of statements (now known as P1–P15) that capture the essential properties of cells, Lusztig’s $\mathbf{a}$-function, the $\gamma$-coefficents and the distinguished involutions. In the case of type $\mathsf{A}$ properties P1–P15 can be proved using the positivity properties of the Kazhdan-Lusztig basis (see [21, §15]). We record some of these properties for later use:

### v2

@@ Introduction
For general affine (equal parameter) type, Lusztig gave a conjectural description of the asymptotic algebra in terms of the Langlands dual group [23]. In the case of affine type $\mathsf{A}$, this conjecture is equivalent to show that, for all two-sided cells $\Delta_{\lambda}$, the subalgebra $\mathcal{J}_{\lambda}=\mathcal{J}_{\Delta_{\lambda}}$ is isomorphic to the representation ring of a certain connected reductive group $F_{\lambda}$ with Weyl group $G_{\lambda}$. This was first proved by Xi in a remarkable paper [34] in 2002 where he gave the first explicit construction of the asymptotic algebra in affine type $\mathsf{A}$ using the notion of chains and antichains, following work of Shi [31]. Around the same time Bezrukavnikov and Ostrik [4] verified Lusztig’s conjecture in a modified form for general affine type (see also [3, 7]). A further approach to constructing the asymptotic algebra in affine type $\mathsf{A}$ was given recently by Kim and Pylyavskyy [17] using the affine matrix ball construction.
The aim of this paper is to present a new construction of the asymptotic algebra in affine type $\mathsf{A}$ as a matrix algebra with coefficients in the ring of $G_{\lambda}$-symmetric functions which we believe is interesting for three main reasons: (1) our construction is based on the combinatorial notion of alcove paths and the representation ring of $F_{\lambda}$ appears very naturally via the $G_{\lambda}$-Schur symmetric functions, (2) an asymptotic version of the Plancherel Theorem is proved, which endows the asymptotic algebra with an inner product, and (3) there are indications that our methods are adaptable to other affine types, and moreover to the unequal parameter case. For instance, similar methods have been used by the second and fourth authors to construct the asymptotic algebra associated to the lowest two-sided cell in general affine type (for all parameters; see [14, Section 6]) and to construct the full asymptotic algebra for all affine Hecke algebras associated to rank $2$ root systems (for all parameters; see [14, 15]).
@@ 1. $\bullet$
    We construct a canonical trace functional on the asymptotic algebra for a weighted Hecke algebra of arbitrary type, following Lusztig [24, §20.1(b)] (see Section 5.2), and we prove an asymptotic version of Opdam’s Plancherel Theorem [27] giving a spectral decomposition of this trace functional in affine type $\mathsf{A}$. In particular, we show that the terms in the Plancherel Theorem (in affine type $\mathsf{A}$) are in bijection with the set of two-sided cells, and that the coefficients that appear are linked to Lusztig’s $\mathbf{a}$-function (see Section 5.3).
@@ 3. $\bullet$
We now describe in more details the content of the paper. To understand the general philosophy it is helpful to understand the situation for the lowest two-sided cell $\Delta_{0}$ (the minimal cell with respect to the two-sided order $\leq_{LR}$; in this case our method applies to all affine Weyl groups and all choice of parameters). This cell corresponds to the partition $(1^{n+1})$, and contains the longest element $\mathsf{w}_{0}$ of the finite Weyl group $W$. Let $\Gamma_{0}$ be the right cell that contains $\mathsf{w}_{0}$. Then $\Gamma_{0}\cap\Gamma_{0}^{-1}$ is the set of words $t_{\gamma}\mathsf{w}_{0}$ where $\gamma\in P_{+}$ is a dominant weight and $t_{\gamma}\in\widetilde{W}$ is the translation by $\gamma$. Let $\pi_{0}$ be the principal series representation of $\widetilde{H}$ as defined in [14] for example. Then $\pi_{0}$ has a combinatorial description in terms of positively folded alcoves paths, is bounded by $\ell(\mathsf{w}_{0})$ and recognises $\Delta_{0}$ (see [14, Section 6]). Furthermore, the leading matrix $\mathfrak{c}_{0}(t_{\gamma}\mathsf{w}_{0})$ for $\gamma\in P_{+}$ has a unique non-zero coefficient, equal to $\mathfrak{s}_{\gamma}(\zeta)$ (the Schur function of type $W$). In this case, the Satake Isomorphism provides an isomorphism between ${\bf 1}_{0}\widetilde{H}{\bf 1}_{0}$ and the ring of $W$-symmetric functions (see [18, 26]). In this paper we construct analogues of the above results for all two-sided cells.
@@
Before turning our attention to the asymptotic Plancherel Theorem, let us first recall the situation for a finite dimensional Hecke algebra $H_{0}$ of type $\mathsf{A}_{n}$. In this case the canonical trace on the Hecke algebra decomposes as a linear combination of irreducible characters (indexed by partitions) and the coefficients that appear are normalisations of the generic degrees of $H_{0}$. It turns out that the coefficient associated to $\lambda$ has valuation $2\mathbf{a}(\mathsf{w}_{\lambda^{\prime}})$ (see [10]). There is an analogue of this decomposition for affine Hecke algebras in the form of the Plancherel Theorem [27] which expresses the canonical trace as a sum of integrals over families of representations. In this paper we show that, in affine type $\mathsf{A}$, a similar phenomenon as in the finite case happens: the terms in the Plancherel formula are in bijection with the two-sided cells, and the coefficients have valuation equal to $2\mathbf{a}(\Delta_{\lambda})$ (this behaviour was first observed in affine Hecke algebras of types $\tilde{\mathsf{G}}_{2}$ and $\tilde{\mathsf{C}}_{2}$ by the second and fourth authors in [14, 15]). This leads to an asymptotic version of the Plancherel Theorem, which in turn gives rise to an inner product on the ring $\mathfrak{C}_{\lambda}$ of leading matrices.
@@
In Section 5 we construct a canonical trace on Lusztig’s asymptotic algebra $\mathcal{J}$ (for arbitrary Coxeter type). We then recall the Plancherel Theorem for type $\tilde{\mathsf{A}}_{n}$, following [2], and develop our asymptotic Plancherel Theorem in Theorem 5.7. This allows us to prove that $\pi_{\lambda}$ recognises $\Delta_{\lambda}$ (Theorem 5.10) and that $\mathcal{J}_{\lambda}\cong\mathfrak{C}_{\lambda}$ (Theorem 5.11).
@@ 1.1 The symmetric group and type $\mathsf{A}_{n}$ root system
The support of a root $\alpha\in\Phi$ is $\mathrm{supp}(\alpha)=\{i\mid c_{i}\neq 0\}$, where $\alpha=\sum_{i=1}^{n}c_{i}\alpha_{i}$. For $J\subseteq\{1,\ldots,n\}$ let $\Phi_{J}=\{\alpha\in\Phi\mid\mathrm{supp}(\alpha)\subseteq J\},$ and for $w\in W$ write $\Phi_{J}(w)=\Phi(w)\cap\Phi_{J}$. If $w=uv$ with $u\in W_{J}$ and $v\in{{}^{J}}W$ then $\Phi(u)=\Phi(w)\cap\Phi_{J}$. In particular, we have ${{}^{J}}W=\{v\in W\mid\Phi_{J}(v)=\emptyset\}$ (see [13, Lemma 1.2]).
@@ 1.2 Partitions and tableaux
Let $\leq$ be the dominance partial order on $\mathcal{P}(n+1)$ given by $\mu\leq\lambda$ if and only if $\mu(i)\leq\lambda(i)$ for all $i\geq 1$ (where we set $\lambda_{i}=0$ if $i>r(\lambda)$). Recall that $\mu\leq\lambda$ if and only if $\lambda^{\prime}\leq\mu^{\prime}$.
@@
Then for $j\in J_{\lambda}$ we have $\langle\alpha_{j},\rho_{\lambda}\rangle=1$ (c.f. [13, Lemma 1.4]).
@@ 1.3 The element $\mathsf{u}_{\lambda}$
Let $\mu=\boldsymbol{\mu}(w,\lambda)$. Let $\gamma_{1},\gamma_{2},\ldots,\gamma_{k}$ be the blocks of the $\lambda$-expression for $w$, arranged so that they are in decreasing length. Thus $\mu=(|\gamma_{1}|,\ldots,|\gamma_{k}|)$. Let $\mathbf{t}$ be a (not necessarily column strict) tableau of shape $\mu$ given by entering the elements of $\gamma_{i}$ into the $i$th row of $\mathbf{t}$ in any order. By construction of the $\lambda$-expression, the elements of row $i$ of $\mathbf{t}$ lie in different columns of $\mathbf{t}_{c}(\lambda^{\prime})$, and so by the Dominance Lemma for partitions (see [30, Lemma 2.2.4]) we have $\mu\leq\lambda^{\prime}$ as required. ∎
@@ 1.4 The extended affine Weyl group
For $\alpha\in\Phi$ and $k\in\mathbb{Z}$ let $H_{\alpha,k}=\{x\in V\mid\langle x,\alpha\rangle=k\},$ and let $s_{\alpha,k}(v)=v-(\langle v,\alpha\rangle-k)\alpha$ be the orthogonal reflection in $H_{\alpha,k}$. The group $W_{\mathrm{aff}}$ is a Coxeter group with generators $S=\{s_{0}\}\cup S_{0}$, where $s_{0}=s_{\varphi,1}$, with $\varphi$ the highest root of $\Phi$. The group $\widetilde{W}$ is not a Coxeter group, however we have
$\widetilde{W}=W_{\mathrm{aff}}\rtimes\Sigma\quad\text{where}\quad\Sigma=P/Q\cong\mathbb{Z}/(n+1)\mathbb{Z}.$
We extend the length function $\ell:W_{\mathrm{aff}}\to\mathbb{N}$ to $\widetilde{W}$ by setting $\ell(w\pi)=\ell(w)$ for all $w\in W_{\mathrm{aff}}$ and $\pi\in\Sigma$. Thus $\Sigma=\{w\in\widetilde{W}\mid\ell(w)=0\}$. Each $\pi\in\Sigma$ induces a permutation of the nodes of the extended Dynkin diagram by $\pi s_{i}\pi^{-1}=s_{\pi(i)}$ for $0\leq i\leq n$. Let $\sigma\in\Sigma$ be the element with $\sigma(i)=i+1$ (with indices read cyclically). By a reduced expression for $w\in\widetilde{W}$ we shall mean a decomposition $w=s_{i_{1}}\cdots s_{i_{\ell}}\pi$ with $\ell=\ell(w)$ and $\pi\in\Sigma$.
@@ 1.5 The extended affine Hecke algebra of type $\tilde{\mathsf{A}}_{n}$
The Bernstein relation (see [22, Proposition 3.6]) is
@@ 1.6 Kazhdan-Lusztig theory
Let $\deg:\mathsf{R}\to\mathbb{Z}\cup\{-\infty\}$ denote degree in $\mathsf{q}$. For example $\deg(\mathsf{q}^{-3})=-3$ and $\deg(0)=-\infty$. Since $h_{x,y,z}$ is bar-invariant we have $\deg(h_{x,y,z})\geq 0$ (provided $h_{x,y,z}\neq 0$), and by [19, §7.2] that $\deg(h_{x,y,z})\leq\ell(\mathsf{w}_{0})$ for all $x,y,z\in\widetilde{W}$.
@@
In particular, note that for $w\in W_{J}$ we have $T_{w}C_{\mathsf{w}_{J}}=C_{\mathsf{w}_{J}}T_{w}=\mathsf{q}^{\ell(w)}C_{\mathsf{w}_{J}},$ and hence $C_{\mathsf{w}_{J}}^{2}=\mathsf{q}^{\ell(w_{J})}W_{J}(\mathsf{q}^{-2})C_{\mathsf{w}_{J}}=\mathsf{q}^{-\ell(\mathsf{w}_{J})}W_{J}(\mathsf{q}^{2})C_{\mathsf{w}_{J}}.$
@@
Since $\ell(sw)<\ell(w)$ for all $s\in J$ we have $C_{s}C_{w}=(\mathsf{q}+\mathsf{q}^{-1})C_{w}$ (see [24, Theorem 6.6]), and hence $T_{s}C_{w}=\mathsf{q}C_{w}$. Thus $T_{v}C_{w}=\mathsf{q}^{\ell(v)}C_{w}$ for all $v\in W_{J}$. The result follows by expanding $C_{\mathsf{w}_{J}}$ in the $T_{w}$ basis using (1.3). ∎
@@
Let $\mathcal{D}=\{w\in W\mid\mathbf{a}(w)=-\deg(P_{e,w})\}$ (the set of distinguished involutions). For $d\in\mathcal{D}$ define $n_{d}\in\mathbb{Z}$ by
$P_{e,d}=n_{d}\mathsf{q}^{-\mathbf{a}(d)}+\text{strictly smaller powers of $\mathsf{q}$}.$
In [24], Lusztig gave a series of statements (now known as P1–P15) that capture the essential properties of cells, Lusztig’s $\mathbf{a}$-function, the $\gamma$-coefficents and the distinguished involutions. In the case of type $\mathsf{A}$ properties P1–P15 can be proved using the positivity properties of the Kazhdan-Lusztig basis (see [24, §15]). We record some of these properties for later use:
@@
  2. P3.
    If $x\in\widetilde{W}$ then there exists a unique $d\in\mathcal{D}$ such that $\gamma_{x,x^{-1},d}\neq 0$.
@@
  4. P5.
    If $d\in\mathcal{D}$ and $x\in\widetilde{W}$ with $\gamma_{x,x^{-1},d}\neq 0$ then $\gamma_{x,x^{-1},d}=n_{d}=\pm 1$.